OpenStudy (anonymous):

The teacher of a seventh-grade class flipped a coin 100 times. The students recorded the results in the table shown below: Side of Coin Landed On Heads Tails Number of Flips 56 44 Which of the following best describes the experimental probability of getting heads? The experimental probability is same as the theoretical probability. The experimental probability is 6% lower than the theoretical probability. The experimental probability is 6% higher than the theoretical probability. The experimental probability cannot be concluded from the data in the table.

Mehek (mehek14):

how many times did they get heads

Mehek (mehek14):

and what's the total

OpenStudy (anonymous):

56

OpenStudy (anonymous):

after you find what @Mehek14 wrote, then compare that to the theoretical probability

OpenStudy (anonymous):

and the total is ...

OpenStudy (anonymous):

110

Mehek (mehek14):

56 + 44 = 100

OpenStudy (anonymous):

we already now that 1/2 is the theoretical probability

Mehek (mehek14):

so \(\dfrac{56}{100}=56\)%

OpenStudy (anonymous):

oh i did plus 54 my bad

OpenStudy (anonymous):

ok thank you!!

Mehek (mehek14):

and as @heretohelpalways said 1/2 = what percentage?

Mehek (mehek14):

what is 1/2 of 100%

OpenStudy (anonymous):

so its C

Mehek (mehek14):

yes correct

OpenStudy (anonymous):

thank you! :)

Mehek (mehek14):

yw ^_^